# Robust Finetuning of Vision-Language-Action Robot Policies via Parameter Merging

### v1

@@ 5.2 Modality-specific merging
We show in Section˜6.3 that, somewhat surprisingly, in many settings it suffices to only merge the parameters of the language model backbone.
@@ 5.3 Continual task adaptation
We observe that RETAIN enables finetuned policies to retain the generalist capabilities of the pre-trained policy. As such, we can use RETAIN to sequentially add tasks into a pre-trained checkpoint by iteratively merging finetuned weights into the base model and continuing to finetune from the merged checkpoint (see Fig.˜4). Formally, for a sequence of target tasks $\eta{1},\dots\eta{N}$ we can compute a sequence of adapted RETAIN policies that accumulate new task capabilities as:
@@ 6 Experiments
The goal of our experiments is to evaluate RETAIN’s ability to robustly finetune generalist policies to new tasks, i.e., to broaden the finetuned policy’s ability to generalize to unseen settings in the target task. Concretely, we aim to answer the following questions: (1) Can RETAIN learn a new skill robustly and generalize more broadly to variations of the skill than prior finetuning approaches? (2) What factors influence whether we can effectively merge pre-trained and finetuned policy? (3) Can RETAIN enable continual merging of skills into pre-trained policies by retaining generality on tasks other than the finetuning task?
@@ 6.1 Experimental Setup
Figure 5: We evaluate policy finetuning on two real-world DROID finetuning tasks (left, middle) and three simulated Libero finetuning tasks (right, only one visualized here). In each task, we collect a modest number of demonstrations (5050150150) in a comparatively narrow setting (yellow), but evaluate on a much broader set of variations for the same task (green), including variations to scene, object instances, initial positions, lighting conditions, distractors, and viewpoints. This tests transfer of the generalization ability of the pre-trained policy to the target task.
@@
We evaluate RETAIN in real-world and simulated finetuning settings (see Fig.˜5). For our real-world experiments, we use the DROID robot setup [24], which consists of a 7-DoF Franka robot arm with a wrist-mounted camera and at least one external camera. We design two challenging fine-tuning tasks: wiping the whiteboard with an eraser (which we call whiteboard) and putting the dishes into a drying rack (which we call plates). For both tasks, we collect roughly $50$ and $100$ demonstrations, respectively. All demonstrations are collected in a single environment with minor position variations (see Fig.˜5, top, “ID”), to mirror the variation in typical narrow-data finetuning regimes [69]. We test generalization of the finetuned policies on a much broader set of target environments (Fig.˜5, bottom, “OOD”), which include unseen backgrounds, object instances, and camera views. For our simulated experiments, we use the Libero simulation environment [30]. We finetune policies pre-trained on the Libero-{object, spatial, goal, 90} datasets to three new tasks: pot-on-stove, mugs-on-plates, and items-into-basket. We use $\approx 45$ demonstrations per task, obtained after filtering and preprocessing the 50 demos provided with the Libero simulator, which only contain minor variations to the initial positions of each object, and again test on a much broader distribution of initial positions, backgrounds, and additional distractors (see Fig.˜5, right). More setup details can be found in Section˜A.5 and Section˜A.6.
For both LIBERO and DROID, we evaluate our method and baselines on three different kinds of OOD scenes. To pick the merging coefficient $\alpha$, we use one OOD scene as the “validation” scene, and tune the hyperparameter $\alpha$ for best performance on that validation scene. We only use $\alpha\in\{0.25,0.5,0.75\}$ in DROID. Then, we use the rest of the OOD scenes as the “test” scenes without any hyperparameter tuning.
Pre-trained policies.
We use state-of-the-art pre-trained robot policies for our experiments. Concretely, for our real-world DROID experiments, we use $\pi_{0}$-FAST-DROID [43], the best open-source DROID policy at the time of our experiments as judged by the RoboArena policy ranking [2] [^0] The current strongest policy, $\pi_{0.5}$ [20], was only open-sourced after our experiments. We look forward to testing RETAIN on $\pi_{0.5}$ in the future.. For Libero, we use a $\pi_{0}$ [6] policy fine-tuned on the Libero-{object, spatial, goal, 90} datasets as our pre-trained policy. $\pi_{0}$ is a generalist policy with a flow-based action expert, while $\pi_{0}$-FAST-DROID is an autoregressive transformer based on next-token prediction.
@@
We compare against prior methods, including those that incorporate regularization techniques to reduce overfitting, in the robust finetuning setting. Concretely, we compare our approach, RETAIN, to: “Task-FT”, which finetunes the pre-trained policy only on the target task dataset using behavioral cloning (Eq.˜1) [5, 7]; “Co-FT”, which finetunes on a mix of pre-training and target task data to reduce overfitting [14, 11], “LoRA”, which uses low-rank adaptation [16] during finetuning to retain more of the pre-training capabilities [39, 25]; “Freeze-FT”, which freezes the language model backbone during finetuning and only updates the vision encoder, and, in the case of $\pi_{0}$, the action expert output head, following similar approaches in prior work, e.g., Kim et al. [25], Zhang et al. [69]; Scratch, which learns a policy from scratch on the demonstration dataset instead of finetuning a generalist policy. For more details about the policy classes and implementations, see Section˜A.8. Section˜A.7 details our choice of hyperparameters and tuning process.
@@ 6.2 RETAIN learn new skills robustly and generalize more broadly
We compare RETAIN with the aforementioned baseline methods when finetuning to a new task, and show the performance on the three types of evaluations. Fig.˜6 shows results on two DROID environments, and Fig.˜7 shows the average of three LIBERO environments. The ID and OOD evaluations show how well a method learns to do the new skill: ID evaluations show whether the method learned to fit the demonstration dataset exactly, and OOD evaluations assess whether the method has learned to generalize to the same task exhibiting variations not seen in the finetuning dataset. The OOD (val) scenes show “the upper bound” in performance when $\alpha$ is allowed to be tuned slightly, while the OOD (test) scenes show the performance with no tuning.
@@
In real-world DROID environments (Fig.˜6), all methods perform much better on ID evaluations after finetuning, showing that the policy has adapted to the new task in the exact same context as the demonstration dataset. On ID evaluations, methods that use regularization, such as LoRA and Co-FT, perform slightly worse on whiteboard, likely because they are too constrained to adapt well to the new task. On OOD (test) evaluations, baseline finetuning methods perform significantly worse: while they can complete the new task with $70-80\%$ success rate in the ID setting, they have $30-50\%$ success rate on average in the OOD setting. This shows that the baseline methods are very sensitive to small variations (such as object change, location change, scene change) and cannot generalize to perform well. In comparison, both RETAIN-task-FT and RETAIN-co-FT perform much better, achieving more than $60\%$ on plates and close to $80\%$ on whiteboard OOD evaluation. Note that this is similar to the policy’s ID evaluation performance on whiteboard, suggesting that RETAIN can perform a generalized new skill with the same performance regardless of variations. Comparing the performance of OOD val and test sets, we see that the performance is sometimes not impacted (whiteboard) and sometimes better with slight hyperparameter tuning (plates). This shows that RETAIN is somewhat robust to hyperparameter, but can get even better when tuned for the particular OOD scene. Overall, in OOD evaluations, RETAIN enables the policy to outperform both the base model and other finetuned models.
@@
Figure 7: RETAIN results averaged over the three LIBERO tasks. Similar trend as Fig.˜6.
@@
In LIBERO environments (Fig.˜7), all methods exhibit trends similar to those in DROID environments. Since the LIBERO simulation is a much easier task than real world robotic tasks, certain baseline fine-tuning methods achieve near-perfect performance in ID evaluations. Under OOD evaluations, both RETAIN-task-FT and RETAIN-co-FT help improve the policy’s robustness to scene variations. We observe that the improvement in OOD performance over baselines such as Co-FT in LIBERO is smaller than that in DROID, and we attribute this to the lack of generalist capabilities of the base model. For DROID, the base model $\pi_{0}\text{-FAST}$ was trained on $76k$ diverse trajectories in $564$ scenes, while the LIBERO base model is only trained on $5.3k$ trajectories in $117$ scenes with fairly limited diversity. As such, the LIBERO base model contains much less generalist ability than the DROID base model, and so merging with it inherits less generalization power, giving less improvement under OOD evaluations. There is also a much bigger gap between val and test OOD scenes, because the different types of variations for each OOD scene has a different difficulty level for the limited generalist model in LIBERO.
@@
RETAIN Still Perform Well on Tasks from the Pretraining Distribution.
@@
As we have shown above in Section˜6.2, RETAIN allows the model to not overfit to the finetuning dataset and generalize to solve a broader distribution of the finetuning task. One natural question is whether RETAIN has overfit to the finetuning task distribution, and whether it can still solve tasks under the pretraining distribution. To evaluate retention of generalist skills, we evaluate RETAIN and baselines under our generalist evaluation scenes. For DROID, we evaluate on $44$ different real-world tasks; for LIBERO, we evaluate on $20$ random tasks in the LIBERO pretraining dataset ($5$ each from Libero-{object, spatial, goal, 90}). Fig.˜6 and Fig.˜7 (second subfig) show that RETAIN performs just as well as the pretrained model on generalist evaluations, showing that the merged model has not lost its ability to solve old tasks seen during pretraining. In particular, RETAIN works even better when combined with co-finetuning; we will offer more discussion on this in Section˜6.2.
@@
In Fig.˜6 and Fig.˜7, RETAIN-co-FT outperforms RETAIN-task-FT in all three evaluation settings in almost all tasks. In particular, RETAIN-co-FT is particularly effective at improving performance in generalist and OOD evaluation. In fact, we observe that co-FT almost always help improve performance over task-FT in generalist evaluation, but not always in OOD evaluation. We hypothesize this is because co-finetuning and model merging play different roles in the regularization process: co-finetuning helps the finetuned model not overfit to the small target dataset by continuous training on pretraining data, but does not help elicit pretrained knowledge to help generalization on the new task; on the other hand, model merging explicitly tries to elicit pretrained knowledge and combine it with finetuning knowledge in parameter space, but doing so with a task-FT model is worse at keeping pretraining abilities because the task-FT model is overfitted to the target dataset.
6.3 Analyzing importance of merging different parameters in RETAIN
Figure 8: RETAIN performs best when most of the weights come from the finetuned model. Plot shows OOD performance on three types of OOD variations averaged across three LIBERO tasks.
@@ 6.3 Analyzing importance of merging different parameters in RETAIN
In this section, we seek to build a mechanistic understanding of how the merging coefficient impacts the performance of merged models. To begin with, we try to understand how changing the coefficient $\alpha$ in Eq.˜2, when all parameters are interpolated with the same value, impacts performance. In Fig.˜8, we plot OOD performance against $\alpha$ averaged across three LIBERO tasks, each with three types of OOD scenes. $\alpha=0$ corresponds to the pretrained model, and $\alpha=1$ corresponds to the co-finetuned model without merging. Model merging ($0<\alpha<1$) helps improve model performance in OOD evaluation, as long as the merged model is not too deviated from the finetuned model. When the merged model is too similar to the pretrained model, it does not have enough task-specific knowledge, and has near 0 performance. And as we have shown in Fig.˜7, the model with the $\alpha$ value that has the best OOD performance is also comparable to baselines in ID evaluations and better at generalist evaluations.
@@
Next, we explore whether merging different parameters with different coefficients has an impact on the merged policy’s performance. Specifically, we consider modality-specific merging in the context of VLA policies. As explained in Section˜5.2, we use separate coefficients for merging the parameters of the vision encoder, language model, and action expert. Fig.˜9 (left) shows the OOD performance of the merged model as we vary $\alpha_{v}$, $\alpha_{l}$, and $\alpha_{a}$ from 0 to $1$ on the mugs-on-plates task: dark colors represent low OOD performance, and light colors represent high OOD performance. Observe that the cube has the largest color gradient in the $\alpha_{l}$ direction: this shows that the language model parameters have the most influence on performance. Interestingly, we see that $\alpha_{l}=1$ does not yield the best performance; the best performing models has $\alpha_{l}=0.8$ (see the highlighted plane at $\alpha_{l}=0.8$ in Fig.˜9 (left) for the brightest colored dots). Next, to understand how $\alpha_{v}$ and $\alpha_{a}$ impact performance, we plot in Fig.˜9 (middle) the change in OOD performance with these two coefficients when averaged over $\alpha_{l}$. This 2D plot essentially squashes the 3D cube plot in the language direction. Unlike the behavior of $\alpha_{l}$, higher values of $\alpha_{a}$ and $\alpha_{v}$ lead to better performance; the best performance is achieved at $\alpha_{a}=\alpha_{v}=1$.
@@
These results suggest that during model merging, it may suffice to only merge the parameters of the language model backbone ($\alpha_{l}<1$) and leave the parameters in the vision encoder and the action expert set to the parameter values in the finetuned model ($\alpha_{a}=\alpha_{v}=1$). To validate this hypothesis, we compare the OOD performance of merging all parameters with RETAIN ($0<\alpha<1$) to only merging language model parameters. In Fig.˜9 (right), we plot the performance of the two merging schemes over three different LIBERO tasks, each averaged over three types of OOD scenes. Somewhat surprisingly, the result shows that the two merging schemes achieve very similar performance, indicating that we only need to merge parameters from the language model backbone (instead of all parameters) to inherit the model’s generalization ability and robustness to variations in the target scene. To the best of our knowledge, this is the first work to demonstrate the importance of different parameter groups for model merging in VLAs. We believe this may inform future work on finetuning VLAs, providing insight on which parameter groups are most critical.
6.4 RETAIN enables learning multiple skills robustly

### v2

@@ 5.2 Modality-Specific Merging
We show in Section˜6.4 that, somewhat surprisingly, it often suffices to only merge the language model parameters.
@@ 5.3 Continual Task Adaptation
We observe that RETAIN enables finetuned policies to retain the generalist capabilities of the pretrained policy. As such, we can use RETAIN to sequentially add tasks into a pretrained checkpoint by iteratively merging finetuned weights into the base model and continuing to finetune from the merged checkpoint (see Fig.˜5). Formally, for a sequence of target tasks $T_{\eta_{1}},\dots T_{\eta_{N}}$ we can compute a sequence of adapted RETAIN policies that accumulate new task capabilities as:
@@ 6 Experiments
The goal of our experiments is to evaluate RETAIN’s ability to robustly finetune generalist policies to new tasks, i.e., to broaden the finetuned policy’s ability to generalize to unseen settings in the target task. Concretely, we aim to answer the following questions: (1) Can RETAIN learn a new skill robustly and generalize more broadly to variations of the skill than prior finetuning approaches? (2) What factors influence whether we can effectively merge pretrained and finetuned policy? (3) Can RETAIN enable continual merging of a sequence of several skills into the pretrained policies?
@@ 6.1 Experimental Setup
Figure 6: We evaluate policy finetuning on two real-world DROID finetuning tasks (left, middle) and three simulated LIBERO finetuning tasks (right, only one visualized here). In each task, we collect a modest number of demonstrations (5050100100) in a comparatively narrow setting (blue), but evaluate on a much broader set of variations for the same task (yellow), including variations to scene, object instances, initial positions, lighting conditions, distractors, and viewpoints. This tests transfer of the generalization ability of the pretrained policy to the target task. Example trajectories in Fig.˜3.
@@
We evaluate RETAIN in real-world and simulated finetuning settings (see Fig.˜6). For our real-world experiments, we use the DROID robot setup [25], which consists of a 7-DoF Franka robot arm with a wrist-mounted camera and at least one external camera. We design two challenging fine-tuning tasks: wiping the whiteboard with an eraser (which we call whiteboard) and putting the dishes into a drying rack (which we call plates). For both tasks, we collect roughly $50$ and $100$ demonstrations, respectively. All demonstrations are collected in a single environment with minor position variations (see Fig.˜6, blue), to mirror the variation in typical narrow-data finetuning regimes [71]. We test generalization of the finetuned policies on a much broader set of target environments (Fig.˜6, yellow), which include unseen backgrounds, object instances, and camera views. See example task trajectories in Fig.˜3. For our simulated experiments, we use the LIBERO simulation environment [31]. We finetune policies pretrained on the LIBERO-{object, spatial, goal, 90} datasets to three new tasks from the LIBERO-10 suite: pot-on-stove, mugs-on-plates, and items-into-basket. We use $\approx 45$ demonstrations per task, obtained after filtering and preprocessing the 50 demos provided with the LIBERO simulator, which only contain minor variations to the initial positions of each object, and again test on a much broader distribution of initial positions, backgrounds, and additional distractors (see Fig.˜6, right). More setup details can be found in Section˜A.5 and Section˜A.6.
For both LIBERO and DROID, we evaluate our method and baselines on three different OOD scenes. To pick the merging coefficient $\alpha$, we use one OOD scene as the “validation” scene, and tune the hyperparameter $\alpha$ for best performance on that validation scene. We only use $\alpha\in\{0.25,0.5,0.75\}$ in DROID. Then, we use the rest of the OOD scenes as the “test” scenes without any hyperparameter tuning.
Pretrained policies.
We use state-of-the-art pretrained robot policies for our experiments. For our real-world DROID experiments, we use $\pi_{0}$-FAST-DROID [45], the best open-source DROID policy at the time of our experiments as judged by the RoboArena policy ranking [3] [^0] The current strongest policy, $\pi_{0.5}$ [21], was only open-sourced after our experiments. We look forward to testing RETAIN on $\pi_{0.5}$ in the future.. For LIBERO, we use a $\pi_{0}$ [7] policy fine-tuned on the LIBERO-{object, spatial, goal, 90} datasets as our pretrained policy. $\pi_{0}$ is a generalist policy with a flow-based action expert, while $\pi_{0}$-FAST-DROID is an autoregressive transformer based on next-token prediction.
@@
We compare against prior methods, including those that incorporate regularization techniques to reduce overfitting, in the robust finetuning setting. Specifically, we compare our approach, RETAIN, to: “Task-FT”, which finetunes the pretrained policy only on the target task dataset using behavioral cloning (Eq.˜1) [6, 8]; “Co-FT”, which finetunes on a mix of pretraining and target task data to reduce overfitting [15, 12], “LoRA”, which uses low-rank adaptation [17] during finetuning to retain more of the pretraining capabilities [41, 26]; “Freeze-FT”, which freezes the language model backbone during finetuning and only updates the vision encoder, and, in the case of $\pi_{0}$, the action expert output head, following similar approaches in prior work, e.g., Kim et al. [26], Zhang et al. [71]; Scratch, which learns a policy from scratch on the demonstration dataset instead of finetuning a generalist policy. For more details about the policy classes and implementations, see Section˜A.8. Section˜A.7 details our choice of hyperparameters and tuning process.
@@ 6.2 RETAIN Learn New Skills Robustly and Generalize More Broadly
We compare RETAIN with the aforementioned baseline methods when finetuning to a new task, and show the performance on the three types of evaluations. Fig.˜7 shows results on two DROID environments, and Fig.˜8 shows the average of three LIBERO environments. The ID and OOD evaluations show how well a method learns to do the new skill: ID evaluations show whether the method learned to fit the demonstration dataset exactly, and OOD evaluations assess whether the method has learned to generalize to the same task exhibiting variations not seen in the finetuning dataset. The OOD (val) scenes show “the upper bound” in performance when $\alpha$ is allowed to be tuned slightly, while the OOD (test) scenes show the performance with no tuning.
@@
In real-world DROID environments (Fig.˜7), all methods perform much better on ID evaluations after finetuning, showing that the policy has adapted to the new task in the exact same context as the demonstration dataset. On ID evaluations, methods that use regularization, such as LoRA and Co-FT, perform slightly worse on whiteboard, likely because they are too constrained to adapt well to the new task. On OOD (test) evaluations, baseline finetuning methods perform significantly worse: while they can complete the new task with $70-80\%$ success rate in the ID setting, they have $30-50\%$ success rate on average in the OOD setting. This shows that the baseline methods are very sensitive to small variations (such as object change, location change, scene change) and cannot generalize to perform well. In comparison, both RETAIN-task-FT and RETAIN-co-FT perform much better, achieving more than $60\%$ on plates and close to $80\%$ on whiteboard OOD evaluation. Note that this is similar to the policy’s ID evaluation performance on whiteboard, suggesting that RETAIN can perform a generalized new skill with the same performance regardless of variations. Comparing the performance of OOD val and test sets, we see that the performance is sometimes not impacted (whiteboard) and sometimes better with slight hyperparameter tuning (plates). This shows that RETAIN is somewhat robust to hyperparameter, but can get even better when tuned for the particular OOD scene. Overall, in OOD evaluations, RETAIN enables the policy to outperform both the base model and other finetuned models.
@@
Figure 8: RETAIN results averaged over the three LIBERO tasks. Similar trend as Fig.˜7.
@@
In LIBERO environments (Fig.˜8), all methods exhibit trends similar to those in DROID environments. Since the LIBERO simulation is a much easier task than real world robotic tasks, certain baseline fine-tuning methods achieve near-perfect performance in ID evaluations. Under OOD evaluations, both RETAIN-task-FT and RETAIN-co-FT help improve the policy’s robustness to scene variations. We observe that the improvement in OOD performance over baselines such as Co-FT in LIBERO is smaller than that in DROID, and we attribute this to the lack of generalist capabilities of the base model. For DROID, the base model $\pi_{0}\text{-FAST}$ was trained on $76k$ diverse trajectories in $564$ scenes, while the LIBERO base model is only trained on $5.3k$ trajectories in $117$ scenes with fairly limited diversity. As such, the LIBERO base model contains much less generalist ability than the DROID base model. And as we will shown in Section˜6.3, merging with a less general base model inherits less generalization power, giving less improvement under OOD evaluations. There is also a much bigger gap between val and test OOD scenes, because the different types of variations for each OOD scene has a different difficulty level for the limited generalist model in LIBERO.
@@
RETAIN still performs well on tasks from the pretraining distribution.
@@
As shown above in Section˜6.2, RETAIN allows the model to not overfit to the finetuning dataset and generalize to solve a broader distribution of the finetuning task. One natural question is whether RETAIN has overfit to the finetuning task distribution, and whether it can still solve tasks under the pretraining distribution. To evaluate retention of generalist skills, we evaluate RETAIN and baselines under our generalist evaluation scenes. For DROID, we evaluate on $44$ different real-world tasks; for LIBERO, we evaluate on $20$ random tasks in the LIBERO pretraining dataset ($5$ each from LIBERO-{object, spatial, goal, 90}). Fig.˜7 and Fig.˜8 (second subfig) show that RETAIN performs just as well as the pretrained model on generalist evaluations, showing that it has not lost its ability to solve old tasks from pretraining. RETAIN works even better when combined with co-finetuning, as discussed below.
@@
In Fig.˜7 and Fig.˜8, RETAIN-co-FT outperforms RETAIN-task-FT in all three evaluation settings in almost all tasks. RETAIN-co-FT is particularly effective at improving performance in generalist and OOD evaluation. In fact, we observe that co-FT almost always helps improve performance over task-FT in generalist evaluation, but not always in OOD evaluation. We hypothesize this is because co-finetuning and model merging play different roles in the regularization process: co-finetuning helps the finetuned model not overfit to the small target dataset by continuous training on pretraining data, but does not elicit pretrained knowledge to help generalization on the new task; on the other hand, model merging explicitly tries to elicit pretrained knowledge and combine it with finetuning knowledge in parameter space; however, doing so with a task-FT model is worse at keeping pretraining abilities because the task-FT model has overfitted to the target dataset.
6.3 RETAIN Scales with the Amount of Pretraining Data
Figure 9: RETAIN performs better on OOD tasks when the pretrained generalist policy is trained with more data. DROID-Subset is trained with $20k$ episodes, DROID-All is all $75k$ episodes of DROID, while DROID-All+PI is trained with all of DROID plus even bigger datasets from Phyiscal Intelligence (PI) [45]. OOD performance is averaged across three scenes on plates.
As we saw in Section˜6.2, RETAIN is not only able to perform well on tasks from the pretraining distribution, but also solve broader variations of finetuning tasks, as if it “merged” the generalist ability of the pretrained model with the task specialization of the finetuned model. This begs the question: how well does RETAIN do when the pretrained model contains differing amounts of “generalist knowledge”, as induced by different amounts of pretraining data? To study how differing amounts of pretraining data affects RETAIN’s performance, we use the DROID dataset, containing a large quantity of diverse robot data. We consider three pretrained generalist VLA policies: (1) the public $\pi_{0}$-FAST-DROID checkpoint that is trained on all of DROID and a large repertoire of robot data from Physical Intelligence [45], (2) only trained on all of DROID, totaling $76k$ episodes, and (3) only trained on the subset of DROID collected at Berkeley and Stanford, totaling $20k$ episodes. All three policies are pretrained by taking roughly $1$ epoch over the dataset, and finetuned on the plates task with the same hyperparameters. We then evaluate the three RETAIN-co-FT policies, each obtained by merging the respective pretrained policy with its finetuned counterpart, on ID and OOD scenes in Fig.˜9. In ID evaluations, DROID-All + PI performs the best, slightly outperforming the two other models pretrained with less data. In OOD evaluations, more pretraining data leads to significantly better performance, with the best model performing nearly as good on OOD evaluation as it does on ID evaluation. This shows that RETAIN scales with the amount of pretraining data, and can better “transfer” generalist knowledge from the pretrained model to new scenes when the pretrained model is more general.
6.4 Analyzing the Importance of Merging Different Parameters in RETAIN
Figure 10: Left: Illustration of three different kinds of OOD scenes we consider in LIBERO, shown in pot-on-stove. Right: Plot shows how the OOD performance on three types of OOD variations changes with the merging parameter $\alpha$, averaged across three LIBERO environments.
@@ 6.4 Analyzing the Importance of Merging Different Parameters in RETAIN
In this section, we seek to build a mechanistic understanding of how the merging coefficient impacts the performance of merged models. To begin, we try understanding how changing the coefficient $\alpha$ in Eq.˜2, when all parameters are interpolated with the same value, impacts performance. In Fig.˜10, we plot OOD performance against $\alpha$ averaged across three LIBERO tasks, each with three types of OOD scenes. $\alpha=0$ corresponds to the pretrained model, and $\alpha=1$ corresponds to the co-finetuned model without merging. Model merging ($0<\alpha<1$) helps improve model performance in OOD evaluation, as long as the merged model is not too deviated from the finetuned model. When the merged model is too similar to the pretrained model, it does not have enough task-specific knowledge, and has near 0 performance. And as we have shown in Fig.˜8, the model with the $\alpha$ value that has the best OOD performance is also comparable to baselines in ID evaluations and better at generalist evaluations.
@@
Next, we explore whether merging different parameters with different coefficients has an impact on the merged policy’s performance. Specifically, we consider modality-specific merging in the context of VLA policies. As explained in Section˜5.2, we use separate coefficients for merging the parameters of the vision encoder, language model, and action expert. Fig.˜11 (left) shows the OOD performance of the merged model as we vary $\alpha_{v}$, $\alpha_{l}$, and $\alpha_{a}$ from 0 to $1$ on the mugs-on-plates task: dark colors represent low OOD performance, and light colors represent high OOD performance. Observe that the cube has the largest color gradient in the $\alpha_{l}$ direction: this shows that the language model parameters have the most influence on performance. Interestingly, we see that $\alpha_{l}=1$ does not yield the best performance; the best performing models has $\alpha_{l}=0.8$ (see the highlighted plane at $\alpha_{l}=0.8$ in Fig.˜11 (left) for the brightest colored dots). Next, to understand how $\alpha_{v}$ and $\alpha_{a}$ impact performance, we plot in Fig.˜11 (middle) the change in OOD performance with these two coefficients when averaged over $\alpha_{l}$. This 2D plot essentially squashes the 3D cube plot in the language direction. Unlike the behavior of $\alpha_{l}$, higher values of $\alpha_{a}$ and $\alpha_{v}$ lead to better performance; the best performance is achieved at $\alpha_{a}=\alpha_{v}=1$.
These results suggest that during model merging, it may suffice to only merge the parameters of the language model backbone ($\alpha_{l}<1$) and leave the parameters in the vision encoder and the action expert set to the parameter values in the finetuned model ($\alpha_{a}=\alpha_{v}=1$). To validate this hypothesis, we compare the OOD performance of merging all parameters with RETAIN ($0<\alpha<1$) to only merging language model parameters. In Fig.˜11 (right), we plot the performance of the two merging schemes over three different LIBERO tasks, each averaged over three types of OOD scenes. Somewhat surprisingly, the result shows that the two merging schemes achieve very similar performance, indicating that we only need to merge parameters from the language model backbone (instead of all parameters) to inherit the model’s generalization ability and robustness to variations in the target scene. To the best of our knowledge, this is the first work to demonstrate the importance of different parameter groups for model merging in VLAs. We believe this may inform future work on finetuning VLAs, providing insight on which parameter groups are most critical.
@@
6.5 RETAIN Enables Robust Learning of Multiple Skills Sequentially